# New GPU developments in the Madgraph CUDACPP plugin: kernel splitting, helicity streams, cuBLAS color sums

### v2

@@ 4.4 Studies of a $2\!\rightarrow\!6$ process: $gg\!\rightarrow\!t\bar{t}gggg$
Table 7: Profiling of Feynman diagrams and color sums for $gg\!\rightarrow\!t\bar{t}gggg$ on GPU in different DCDIAG modes and with different numbers of diagrams per group. CUDA/C++ code generated using a modified version of CUDACPP, instrumented with rdtsc timers (hack_ihel6p2_ggtt4g branch). Results for an NVidia A100 GPU. The tests are executed with CUDACPP_RUNTIME_GOODHELICITIES=ALL to bypass helicity filtering. The average ME value shown in the table, computed from the same random numbers in all cases, is used as a basic validation of physics results.
For what concerns the GPU implementation, diagram splitting is also a game changer, allowing the first (to my knowledge) CUDACPP calculation of matrix elements on GPUs for $gg\!\rightarrow\!t\bar{t}gggg$. As shown in Table 5, CUDA builds succesfully complete in less than 10 minutes, in both modes DCDIAG=0 and DCDIAG=1, if the number of diagrams per group is kept below 100. If dpg is increased to 1000, the DCDIAG=0 build succeeds in 3h, while the DCDIAG=1 build fails because it requires too much memory: it was specifically this point that motivated me to get rid of templates, in an attempt to simplify the task of the CUDA compiler, but unfortunately the situation does not improve after removing templates (ihel6p2). This is unfortunate, because the DCDIAG=1 build clearly gives higher peak throughputs, and reaches them with smaller grid sizes, compared to DCDIAG=0. This is not surprising, because the latter involves much heavier read/write access of wavefunctions and dual amplitudes in GPU global memory, while the former only uses GPU registers in most cases.
@@ 4.4.1 Profiling of color sums in $gg\!\rightarrow\!t\bar{t}gggg$
The results I have presented for $gg\!\rightarrow\!t\bar{t}gggg$ in Tables 4 and 5 refer to the overall throughput for computing MEs, which includes the calculations of both Feynman diagrams and QCD color sums. Since kernel splitting into diagram groups and the two different DCDIAG options are only relevant to the former and not the latter, I found it useful to separately profile these two components. In doing this, I also tested the impact of the various color sum options in $gg\!\rightarrow\!t\bar{t}gggg$. I only studied a limited number of scenarios, including in particular those which had achieved the highest overall throughputs in the tests presented above.
The results for the SIMD implementation, shown in Table 6, show that the choice of dpg only affects the calculation of Feynman diagrams, as expected. The fraction of time spent in color sums is around 10–12% in the dpg scenario with the fastest calculation of Feynman diagrams: this number is somewhat lower than I had expected, but it is still significantly larger than the 5–10% observed for the simpler $gg\!\rightarrow\!t\bar{t}ggg$ process. The SIMD speedups of the color sum alone, which is computed here in single precision (FPTYPE=m), are also lower than I had expected: for instance, a factor around 9 is observed for the 512z implementation, which is closer to the x8 speedup expected in double precision than to the x16 speedup expected in single precision. This may deserve further investigations in the future [^10] From a first quick analysis, the problem may be that, for FPTYPE=m on CPUs, the quadratic form $J^{H}(C)J$ is maybe not fully computed in single precision: the first multiplication $H(C)J$ is computed in single precision, but the second multiplication $J^{H}\left[(C)J\right]$ seems to be computed in double precision. .
In this context, it should also be noted that cuBLAS color sums provide physics results with the same accuracy as those of GPU kernels, except in TF32 math mode. In Tables 6 and 7, in particular, I explicitly included the average ME value of my simple tests: this is the same to seven significant digits in all scenarios I considered both on CPUs and GPUs, when starting from the same random numbers; the only exception is cuBLAS/TF32, where discrepancies of order 10-4 are observed. As discussed in Sec. 3.3, the TF32 option is a switch that I introduced to encourage cuBLAS to use tensor cores. As shown in 7, this does provide a moderate speedup of color sums by approximately 30%, but it seems quite clear that the overall benefit (reducing the color sum fraction from 3% to 2.3%) is not worth the cost of a loss in physics precision.
@@
Table 8: Run time performance, including profiling of Feynman diagrams and color sums, of $gg\!\rightarrow\!t\bar{t}ggggg$ in different SIMD modes (FPTYPE=m). C++ code generated using CUDACPP hack_ihel6p1, with 1000 diagrams per group and per file (this is the $gg\!\rightarrow\!t\bar{t}ggggg$ code described in Table 2). The code was later instrumented with rdtsc timers for color sum profiling. Results for a single core of an Intel Xeon Gold 6326 CPU. The tests are executed with CUDACPP_RUNTIME_GOODHELICITIES=ALL to bypass helicity filtering, which is still performed sequentially one helicity at a time. As a basic validation of physics results, it was only checked that the five calculations return exactly the same average ME value, when starting from the same random number seeds.
@@ 4.5 Studies of a $2\!\rightarrow\!7$ process: $gg\!\rightarrow\!t\bar{t}ggggg$
Taking this one step further, using the diagram splitting in ihel6p1 I also performed a first quick analysis of the $2\!\rightarrow\!7$ process $gg\!\rightarrow\!t\bar{t}ggggg$. Since code generation for this process takes 30h, I only attempted one configuration that I thought could be promising, using 1000 diagrams per group. It turns out that, both for the C++ and CUDA builds, the main challenge in this case came not from Feynman diagrams but from the size of the 5k x 5k color matrix, which in CUDACPP is defined as a constexpr. In the case of C++, I was able to address the issue by modifying the constexpr limit in the compiler: this was enough for the build to succeed (in less than 1h), and I was then able to obtain the first CUDACPP results from C++ for a $2\!\rightarrow\!7$ process, which are shown in Tab. 8. In the case of CUDA, I used a similar trick to successfully complete compilation (replacing constexpr compile-time evaluations by runtime const evaluations), but linking failed with the relocation truncated to fit error, which signals that the executable would be too large for the linker to handle. I tried several options to bypass this issue, but none of them succeeded. Concerning runtime performance of the C++ code, the results in Tab. 8 are somewhat puzzling, as the observed SIMD speedups are much lower than one would expect. Unlike the $gg\!\rightarrow\!t\bar{t}gggg$ calculation, this seems to suggest that CPU memory access may be a limiting factor for this calculation. In the table, I also provide the detailed profiling of the Feynamn diagram and color sum components, derived with the same tools used for simpler processes: the fact that color sums have a much better SIMD speedup than Feynman diagrams, and that their overall footprint is relatively low (around 5 to 15%) is an interesting and somewhat surprising indication that the Feynman diagram calculation is the bottleneck for such a complex calculation.
@@ 5 Summary and outlook
The second large batch of enhancements, which includes three further sets of software changes, was still work in progress at the time of my previous preprint bib:preprintV1 of this paper, but it has now been completed and I also recommend its inclusion in a new production release of the software. It is available as a github tag bib:ihel6tag , and a pull request for its merge upstream has been created. These three sets of changes all consist in different ways to split up the computation of Feynman diagrams into separate GPU kernels. The fourth development, which uses a different GPU kernel for each Feynman diagram, has a very poor performance, even when the workflow is orchestrated by CUDA graphs, because it implies very heavy access to GPU global memory. The fifth development allows the definition of kernels that compute larger groups of Feynman diagrams, possibly defined in separate source code files. The sixth development goes back to a single GPU kernel for all Feynman diagrams, which however delegates their calculation to different device functions for separate diagram groups. All these changes are fully functional and have been extensively tested on all platforms.
While this second line of R&D started off as an attempt to further improve throughputs for our standard candles such as $gg\!\rightarrow\!t\bar{t}ggg$, it now seems clear to me that these processes are best computed using the traditional CUDACPP approach of a single GPU kernel for all Feynman diagrams, since they are simple enough to afford this: the same approach and the same throughputs are still achievable using the new software, which can be configured to produce a single diagram group. The interest of the new feature, conversely, is in much more complex physics calculations such as those required by $2\!\rightarrow\!6$ and $2\!\rightarrow\!7$ processes. I have shown, in particular, that diagram splitting makes it possible to compute matrix elements for $gg\!\rightarrow\!t\bar{t}gggg$ in both the GPU and vector CPU implementations of CUDACPP, and for $gg\!\rightarrow\!t\bar{t}ggggg$ in the latter. The code I used for the $gg\!\rightarrow\!t\bar{t}gggg$ tests that I presented is also available as a github tag bib:ihel6tag4g , which I created only for documentation purposes and not as a patch to integrate upstream.
@@ Acknowledgements
I thank the MG5aMC and the Madgraph on GPU development teams for our fruitful collaboration on these projects. In particular, I warmly thank Olivier Mattelaer for our initial discussions and our coding together for the new developments shown in this paper during the 2022 GPU Hackathon and more generally for our effective collaboration over many years on the Madgraph on GPU developments. This project would not have been possible without him. I thank Daniele Massaro for useful discussions about SIMD support in CUDACPP and for his help in the integration of this work in new production releases of the CUDACPP software. I warmly thank Domenico Giordano for many useful discussions about hardware benchmarking. I am grateful to the organisers and our mentors at the CSCS GPU Hackathon in Lugano in September 2022, and I thank the other participants from our team, Stephan Hageböck, Stefan Roiser and Zenny Wettersten, for our useful discussions. I kindly acknowledge the use of LUMI HPC resources under project 465001592 (CERN / Madgraph GPU porting, EHPC-DEV-2024D11-007) to derive the results I showed for AMD GPUs.

### v3

@@ 4.4 Studies of a $2\!\rightarrow\!6$ process: $gg\!\rightarrow\!t\bar{t}gggg$
Table 8: Profiling of Feynman diagrams and color sums for $gg\!\rightarrow\!t\bar{t}gggg$ on GPU in different DCDIAG modes and with different numbers of diagrams per group (FPTYPE=m). CUDA/C++ code generated using a modified version of CUDACPP, instrumented with rdtsc timers (hack_ihel6p2_ggtt4g branch). Results for an NVidia A100 GPU (using CUDA 12.4 and gcc 11.5). The tests are executed with CUDACPP_RUNTIME_GOODHELICITIES=ALL to bypass helicity filtering. The average ME value shown in the table, computed from the same random numbers in all cases, is used as a basic validation of physics results.
For what concerns the GPU implementation, diagram splitting is also a game changer, allowing the first (to my knowledge) CUDACPP calculation of matrix elements on GPUs for $gg\!\rightarrow\!t\bar{t}gggg$. As shown in Table 6, CUDA builds succesfully complete in less than 10 minutes, in both modes DCDIAG=0 and DCDIAG=1, if the number of diagrams per group is kept below 100. If dpg is increased to 1000, the DCDIAG=0 build succeeds in 3h, while the DCDIAG=1 build fails because it requires too much memory. It was specifically this point that motivated me to get rid of templates, in an attempt to simplify the task of the CUDA compiler, but unfortunately the DCDIAG=1 build continues to fail after removing templates (ihel6p2). This is a pity, because the DCDIAG=1 build clearly gives higher peak throughputs, and reaches them with smaller grid sizes, compared to DCDIAG=0 (this is not too surprising, as the latter involves much heavier read/write access of wavefunctions and dual amplitudes in GPU global memory, while the former only uses GPU registers in most cases).
@@ 4.4.1 Profiling of color sums in $gg\!\rightarrow\!t\bar{t}gggg$
The results I have presented for $gg\!\rightarrow\!t\bar{t}gggg$ in Tables 5 and 6 refer to the overall throughput for computing MEs, which includes the calculations of both Feynman diagrams and QCD color sums. Since kernel splitting into diagram groups and the two different DCDIAG options are only relevant to the former and not the latter, I found it useful to separately profile these two components. In doing this, I also tested the impact of the various color sum options in $gg\!\rightarrow\!t\bar{t}gggg$. I only studied a limited number of scenarios, including in particular those which had achieved the highest overall throughputs in the tests presented above.
@@
Table 9: Run time performance, including profiling of Feynman diagrams and color sums, of $gg\!\rightarrow\!t\bar{t}ggggg$ in different SIMD modes (FPTYPE=m). C++ code generated using CUDACPP hack_ihel6p1, with 1000 diagrams per group and per file (this is the $gg\!\rightarrow\!t\bar{t}ggggg$ code described in Table 2). The code was later instrumented with rdtsc timers for color sum profiling. Results for a single core of an Intel Xeon Gold 6326 CPU (using gcc 11.5). The tests are executed with CUDACPP_RUNTIME_GOODHELICITIES=ALL to bypass helicity filtering, which is still performed sequentially one helicity at a time. The average ME value shown in the table, computed from the same random numbers in all cases, is used as a basic validation of physics results.
The results for the SIMD implementation, shown in Table 7, show that the choice of dpg only affects the calculation of Feynman diagrams, as expected. The fraction of time spent in color sums is around 10–12% in the dpg scenario with the fastest calculation of Feynman diagrams: this number is somewhat lower than I had expected, but it is still significantly larger than the 5–10% observed for the simpler $gg\!\rightarrow\!t\bar{t}ggg$ process. The SIMD speedups of the color sum alone, which is computed here in single precision (FPTYPE=m), are also lower than I had expected: for instance, a factor around 9 is observed for the 512z implementation, whereas a x16 speedup could in principle be expected for this single precision calculation. This observation was essentially the main motivation for the additional tests and optimization of SIMD color sums that I have described in Sec. 3.5. As discussed therein, it is now clear that this lower-than-expected speedup does not indicate a suboptimal implementation of SIMD color sums, but rather the fact that the none build mode does not represent a good no-SIMD reference scenario for this type of comparisons, as the color sum actually benefits from auto-vectorization in that case.
In this context, it should also be noted that cuBLAS color sums provide physics results with the same accuracy as those of GPU kernels, except in TF32 math mode. In Tables 7 and 8, in particular, I explicitly included the average ME value of my simple tests: this is the same to seven significant digits in all scenarios I considered both on CPUs and GPUs, when starting from the same random numbers; the only exception is cuBLAS/TF32, where discrepancies of order 10-4 are observed. As discussed in Sec. 3.3, the TF32 option is a switch that I introduced to encourage cuBLAS to use tensor cores. Table 8 shows that this does provide a moderate speedup of color sums by approximately 30%, but it seems quite clear that the overall benefit (reducing the color sum fraction from 3% to 2.3%) is not worth the cost of a loss in physics precision.
@@ 4.5 Studies of a $2\!\rightarrow\!7$ process: $gg\!\rightarrow\!t\bar{t}ggggg$
Taking this one step further, I used the diagram splitting functionality of ihel6p1 to also perform a first quick analysis of the $2\!\rightarrow\!7$ process $gg\!\rightarrow\!t\bar{t}ggggg$. Since code generation for this process takes 30h, I only attempted one configuration that I thought could be promising, using 1000 diagrams per group. It turns out that, both for the C++ and CUDA builds, the main challenge in this case came not from Feynman diagrams but from the size of the 5k x 5k color matrix, which in CUDACPP is defined as a constexpr. In the case of C++, I was able to address the issue by modifying the constexpr limit in the compiler: this was enough for the build to succeed (in less than 1h), and I was then able to obtain the first CUDACPP results from C++ for a $2\!\rightarrow\!7$ process, which are shown in Tab. 9. In the case of CUDA, I used a similar trick to successfully complete compilation (replacing constexpr compile-time evaluations by runtime const evaluations), but linking failed with the relocation truncated to fit error, which signals that the executable would be too large for the linker to handle. I tried several options to bypass this issue, but none of them succeeded.
Concerning runtime performance of the C++ code, the results in Tab. 9 are somewhat puzzling, as the observed SIMD speedups are much lower than one would expect. Unlike the $gg\!\rightarrow\!t\bar{t}gggg$ calculation, this seems to suggest that CPU memory access may be a limiting factor for this calculation. In the table, I also provide the detailed profiling of the Feynamn diagram and color sum components, derived with the same tools used for simpler processes: the fact that color sums have a much better SIMD speedup than Feynman diagrams, and that their overall footprint is relatively low (around 5 to 15%) is an interesting and somewhat surprising indication that the Feynman diagram calculation is the bottleneck for such a complex calculation.
@@ 5 Summary and outlook
The second large batch of enhancements, which includes three further sets of software changes, was still work in progress at the time of my first preprint bib:preprintV1 of this paper, but it has now been completed and I also recommend its inclusion in a new production release of the software. It is available as a github tag bib:ihel6tag , and a pull request for its merge upstream has been created. These three sets of changes all consist in different ways to split up the computation of Feynman diagrams into separate GPU kernels. The fourth development, which uses a different GPU kernel for each Feynman diagram, has a very poor performance, even when the workflow is orchestrated by CUDA graphs, because it implies very heavy access to GPU global memory. The fifth development allows the definition of kernels that compute larger groups of Feynman diagrams, possibly defined in separate source code files. The sixth development goes back to a single GPU kernel for all Feynman diagrams, which however delegates their calculation to different device functions for separate diagram groups. All these changes are fully functional and have been extensively tested on all platforms.
While this second line of R&D started off as an attempt to further improve throughputs for our standard candles such as $gg\!\rightarrow\!t\bar{t}ggg$, it now seems clear to me that these processes are best computed using the traditional CUDACPP approach of a single GPU kernel and a single diagram group for all Feynman diagrams, since they are simple enough to afford this: the same approach and the same throughputs are still achievable using the new software, which can be configured to produce a single diagram group. The interest of the new feature, conversely, is in much more complex physics calculations such as those required by $2\!\rightarrow\!6$ and $2\!\rightarrow\!7$ processes. I have shown, in particular, that diagram splitting makes it possible to compute matrix elements for $gg\!\rightarrow\!t\bar{t}gggg$ (which involves over 15k Feynman diagrams) in both the GPU and vector CPU implementations of CUDACPP, and for $gg\!\rightarrow\!t\bar{t}ggggg$ (over 230k diagrams) in the latter. The code I used for the $gg\!\rightarrow\!t\bar{t}gggg$ tests that I presented is also available as a github tag bib:ihel6tag4g , which I created only for documentation purposes and not as a patch to integrate upstream. Another github tag bib:csmtag , which instead I recommend to merge upstream, includes a few minor optimizations of SIMD color sums that I developed in the context of these tests of $2\!\rightarrow\!6$ and $2\!\rightarrow\!7$ processes.
@@ Acknowledgements
I thank the MG5aMC and the Madgraph on GPU development teams for our fruitful collaboration on these projects. In particular, I warmly thank Olivier Mattelaer for our initial discussions and our coding together for the new developments shown in this paper during the 2022 GPU Hackathon, and more generally for our effective collaboration over many years on the Madgraph on GPU developments. This project would not have been possible without him. I thank Daniele Massaro for useful discussions about SIMD support in CUDACPP and for his help in the integration of this work in new production releases of the CUDACPP software. I warmly thank Domenico Giordano for many useful discussions about hardware benchmarking. I am grateful to the organisers and our mentors at the CSCS GPU Hackathon in Lugano in September 2022, and I thank the other participants from our team, Stephan Hageböck, Stefan Roiser and Zenny Wettersten, for useful discussions. I kindly acknowledge the use of LUMI HPC resources under project 465001592 (CERN / Madgraph GPU porting, EHPC-DEV-2024D11-007) to derive the results I showed for AMD GPUs.